A voltaic cell is constructed with two Magnesium (Mg) electrodes. The two cell compartments have [Mg] = 0.24 M and [Mgº] =
Answer #1

solution :- Both are ma elechrocles with different Concentrating so it is a concentration Cell. This cell will work ý Concentusing Nernst equation, Bell - Eéen -0.053 log on to = 0-0.059 log o 24 25 5 -0.059 X-0.3447 = 0.0/02 v ? Econ=70.0102) v Sinu